Complete { (2, 0, 3)} to form an orthogonal



basis of R³

1
Expert's answer
2021-12-06T17:15:37-0500

Let us complete "\{ (2, 0, 3)\}" to form an orthogonal basis. Consider the vector "(3,0,-2)." Since "2\cdot 3+0\cdot0+3\cdot(-2)=0," we conclude that the inner product of "(2, 0, 3)" and "(3,0,-2)" is zero, and hence the vectors "(2, 0, 3)" and "(3,0,-2)" are orthogonal. Further, consider the vector "(0,1,0)". It follows that "2\cdot 0+0\cdot 1+3\cdot 0=0" and "3\cdot 0+0\cdot 1-2\cdot 0=0," and hence the vector "(0,1,0)" is orthogonal to the vectors "(2, 0, 3)" and "(3,0,-2)." Therefore, "\{ (2, 0, 3),(3,0,-2),(0,1,0)\}" is an orthogonal basis of "\R^3."